OpenStudy (megannicole51):

Find the sum -16+4-(4/4)+(4/4^2)-(4/4^3)+....-(4/4^19)

OpenStudy (agent0smith):

It looks like the common ratio is -1/4... each term is multiplied by -1/4 to get the next one. So you can just use a geometric series sum (idk if you have to show calculusey proof of it...)

OpenStudy (megannicole51):

yeah ive gt the common ratio but idk how to find the sum....and nah its online homework

OpenStudy (anonymous):

@agent0smith smarticles!

OpenStudy (agent0smith):

Use the sum formula for a geometric series \[\Large S_n = \frac{ a_n (1- r^n) }{ 1-r }\] from memory that's it... i'll check.

OpenStudy (megannicole51):

that looks familiar! lol

OpenStudy (agent0smith):

I was a bit off.... http://www.regentsprep.org/regents/math/algtrig/ATP2/ArithG12.gif \[\Large S_n = \frac{ a_1 (1- r^n) }{ 1-r }\]

OpenStudy (megannicole51):

okay:) now what lol

OpenStudy (megannicole51):

is a1 the first term?

OpenStudy (agent0smith):

a1 is the first term.. which is? r we know is -1/4 n is the number of terms, which we can find

OpenStudy (megannicole51):

is it -16?

OpenStudy (agent0smith):

a1 = -16 yep. r=-1/4 and the last term is -(4/4^19) which we can use to find n

OpenStudy (anonymous):

math teacher skills smh

OpenStudy (megannicole51):

okay:)

OpenStudy (agent0smith):

K so we just need to find n, the number of terms...

OpenStudy (megannicole51):

21 terms?

OpenStudy (agent0smith):

I think so... 21 looks right. Now plug it all into the formula

OpenStudy (megannicole51):

(-16(1-(-1/4)^21))/(1-(-1/4))

OpenStudy (megannicole51):

=-12.80000 yay its right! thank you so much!

OpenStudy (agent0smith):

:)

OpenStudy (megannicole51):

can you help me with one more? ill post it on a different thing!